Systems Staff Meeting Notes 12/7/98

Systems Staff Meeting Notes
December 7, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng

Announcements
  1. The new Christmas tree is here!
  2. Winston gave Erika in Personnel his choice for the Programmer position. Erika will need to verify the references and will get back to Winston within a day or two.
  3. Patrick Kyle from Telecommunications gave Bill the administrative training for the phones.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We've received a shipment of hard drives in the wrong size: came as 3.5 Gig, should be 6.4 Gig.
  2. When using Outlook to assign a task with an attachment, sending it back & forth will cause Outlook to get hung up. This may have to do with not having the proper software needed to open the attachment.
  3. The ColorLaser Printer in Preservation needs a new hard drive. It is covered under AppleCare.
  4. The phone in Sharon's old office is dead, need to check on the connections.
Liaisons
  1. Need to put out the word for the Liaisons regarding how to handle after-hours problems. We will get the appropriate information up on the Systems web page.
Networking
  1. New CORE POP was installed for Criser Sunday morning. The CSE router has been responsible for 30-40% of campus and they are trying to more evenly distribute that task. We received a notice on Thursday that they were planning to move the Business subnets on to it as well. When the installation occurred, Library East went offline. We had no email or printing capabilities. LAC went down.
    The problem was that they were unable to get to the port that feeds into this building. The fusion splice failed. The problem was not resolved until midnight.
    We will move subnets 238, 239, 54 & 223 to hook up to the Criser POP eventually - maybe even this weekend.
    USI did the fiber work, they had to borrow equipment from GRU. We can expect more reconfiguration. Bill will speak with Dave Pokorney at NERDC regarding the ramifications of the East outage.
  2. MDI SCSI Express on the CD server had a fainting spell. A reboot fixed it on Saturday.
  3. Ashley installed the new switch in Resource Services last Monday night.
  4. New mail server - when we were first trying to bring it up, it refused to be added to the site while the other mail server was in service. After prolonged telephone correspondence with Microsoft, it was determined that a certain number within the parameters had to be replaced: 666! Once the mark of the beast was removed, the server came up.
  5. A kiosk is being installed in the foyer (between the two front doors) of Library West. We gave Al Marrin an IP from our server. Electricity, mounting & wiring had to be done to accommodate this location. It's plugged into a hub, not a specific port on the switch. Used work counter #6, PD 16 - #6 in the closet in room 219 West.
  6. We have had two different reports of hacker activity in the Libraries recently. With the frequency that such activity has been occurring, there may be need for a full-time network security person on campus.
  7. The FCLA FTP site was just hacked & people were uploading games, clogging the system.
Software Projects Status
  1. Winston is trying to get Mellon to work as stand-alone, without using the Borland database engine. The filtering is not working.
  2. Brian has been working on the newspaper conversion project. Searching by institution & publication is close to being ready. The requirements have been knocked down to critical needs.
  3. Personnel - they are down to the point where they only need to convert their Mac files to DOS. They should be off Macs by the end of the month.
DP Coordinator
  1. SPOT tapes were copied by NERDC and delivered to Maps this morning.
  2. Sharon Wheeler received the information on how to reformat the tapes from Compustat to make them readable. She has not started the process yet.
  3. It was determined that the SPOT tape conversion to CDs that was taking place in collaboration with the Geography department was actually in violation of the SPOT licensing policy. We will try to get permission from SPOT to either convert the tapes we have to CD or buy them ready-made in CD format.
Calendar
  1. Suzy out on leave this week, December 7-11.
  2. Ashley out December 21 & 22.
  3. Rich out December 21-28, back December 29.
  4. Debra out December 22-28, back December 29.
  5. Bill out December 21-31, back January 4.
  6. Bobbie out December 24.
  7. Winston out December 29.

Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu